1. Which pair of numbers has 5 in places in which the value of one of the 5s is 10 times greater than in the other number?
A. 105 and 501
B. 6,450 and 3,952
C. 785 and 253
D. 534 and 195
2. At birth, a human baby has a total of 270 bones. This number decreases to 207 in some adults. How many times greater is the value of the 7 in 270 than in 207?

Found 2 solutions by richwmiller, greenestamps:
Answer by richwmiller(17219)

One question at a time
Glad to help you learn the rules and be more considerate of others.

Answer by greenestamps(13195)



1. None of them.

"10 times greater than x" means x, plus x 10 more times -- i.e., 11x. The place values of the 5's in the given numbers are either 5, 50, or 500; none of those numbers is 11 times another of them.

If you mean "10 times AS GREAT", then again the place values are either 5, 50, or 500. Find one pair where the values of the two 5s are 5 and 50, or 50 and 500.

2. The value of the "7" in 270 is 7*10 = 70. The value of the "7" in 207 is 7*1 = 7. 70 is 9 times greater than 7. (70 = 7+63 = 7 + 9 times 7, so 9 times GREATER THAN 7). 70 is clearly 10 TIMES AS GREAT AS 7; it's not 10 times GREATER THAN 7.